# Possible Causes of False General Relativity Violations in Gravitational Wave Observations

### v1

@@ 2.4 Gaps in the Data
Not much work has gone yet to study the effect of data gaps in XG detectors, but some work already exists for data gaps in space-based detectors, from which we can extrapolate some conclusions. Previous work has shown that data gaps can deteriorate and bias parameter estimation for certain sources (Carre & Porter, 2010; Dey et al., 2021), in particular when the data gap coincides with the merger phase. In general, we would expect that a data gap during the merger would inhibit our ability to constrain deviations from GR at high PN order, while gaps in the early inspiral will be the same for low (or negative) PN order modifications to GR. In particular, if the data has a gap, but our analysis does not account for it, parameter correlations between non-GR and GR parameters are likely to introduce biases that may lead to a false GR violation. Certain methods, such as Bayesian data augmentation (Baghi et al., 2019), however, can be used to include missing data periods as auxiliary variables when sampling the posterior distribution of model parameters that have shown promise at eliminating biases.
@@ 2.5 Detector Calibration Error
The GW strain data $d$ is not directly recorded by the interferometer. Instead, it is reconstructed from the voltage $v(f)$ measured by photodetectors and a response function $R(f)$ that relates the digital readout and GW strain, i.e., $d(f)=R(f)v(f)$ (Abbott et al., 2017b). The calibration process includes a series of measurements to construct a reference model for the response function (Abbott et al., 2017b; Tuyenbayev et al., 2017; Viets et al., 2018). Bias in any step of this process can lead to errors in the measured strain data, and systematic errors in parameter estimation could arise if the calibration error is not accounted for. Vitale et al. (2012) investigate the consequences of calibration error in Bayesian inference of source parameters. They find that parameters that suffer the largest biases are those mostly related to the amplitude of GW signals. This implies that calibration errors could have a minor effect in parameterized tests of GR that modify the phase of waveform. They also conclude that $<20$% of amplitude calibration error or $<10-20^{\circ}$ of phase calibration error should not lead to significant biases for all but the strongest signals in the advanced LIGO era, consistent with Payne et al. (2020) and Hall et al. (2019). However, such level of calibration systematics may not be tolerable in the XG era where SNR values could shoot up to hundreds or to even thousands (Read, 2023), since the statistical error scales as 1/SNR while systematics like calibration errors do not. Therefore, it is crucial to improve the calibration techniques along with the sensitivity in the XG era (Essick, 2022; Capote et al., 2024).
@@
It is possible to quantify and mitigate calibration errors in detection and data analysis. The uncertainty of the response function can be indicated by the photon calibrators which apply a known radiation pressure directly on the test masses within the detector (Goetz et al., 2009; Karki et al., 2016; Bhattacharjee et al., 2021; Abbott et al., 2017b). Abbott et al. (2016d) reported $<10$% calibration uncertainty in the strain amplitude and $<5^{\circ}$ in phase during the first observing run of LIGO-Virgo, and in the third observing run these uncertainties were reduced to $<7$% and $<4^{\circ}$, respectively (Sun et al., 2021). Note that these are overall uncertainties, and systematic errors alone are even smaller. These estimates on calibration uncertainties are used as priors to marginalize uncertainties in the GW strains during parameter estimation, which effectively mitigates the calibration error (Farr et al., 2014; Vitale et al., 2021). However, this technique might conceal tiny deviations from GR, since it marginalizes over some level of uncertainties on amplitude and phase. Hence the effect of calibration errors on tests of GR needs to be studied for current and future GW detectors, so that it can be ruled out (or included) as one of the possible causes for false GR violations.
@@ 3.1 Missing Physics in Waveform Models
The current state-of-the-art waveform models used in tests of GR still lack certain physical effects, such as eccentricity of the binary’s orbit, overtones and non-linearities in the ringdown phase of the binary merger, etc. Including each of these known physical effects individually is crucial for precision GR tests, but their collective inclusion is essential for unbiased assessments of GR. Here we discuss missing physical effects that could lead to a false GR violation.
@@ 3.1.1 Eccentricity
The eccentricity of a binary’s orbit depends on the formation history of the binary. Binaries formed through isolated formation channels in the galactic field are expected to have negligible eccentricity when observed in the frequency band of ground-based detectors, whereas binaries inside dense stellar environments such as globular clusters and nuclear star clusters might have moderate to high eccentricities when observed by these detectors. In an isolated formation channel (Mapelli, 2021), the binary goes through various mass transfer episodes between its components, and as the components evolve and undergo supernova explosions, the binary orbit could gain some eccentricity due to supernova kicks. However, due to the emission of gravitational radiation (Peters, 1964; Tucker & Will, 2021) the binary’s orbit shrinks, and the binary sheds away all its eccentricity over the long inspiral, leaving it with negligible eccentricity close to merger (Peters, 1964). For example, if a binary with an initial orbital eccentricity of $0.2$ emits GWs whose dominant mode has a frequency of $0.1$ Hz, the eccentricity reduces to $\sim 10^{-3}$ when it reaches a dominant mode GW frequency of $10$ Hz. That is why binaries detected by LIGO/Virgo are expected to be quasi-circular. On the other hand, a fraction of dynamically formed binaries can still have some eccentricity (and as high as $\sim 1$ at $10$ Hz) when observed in the frequency band of the LIGO/Virgo detectors (Wen, 2003; O’Leary et al., 2009; Bae et al., 2014; Antonini et al., 2016; Silsbee & Tremaine, 2017; Samsing, 2018; Rodriguez et al., 2018; Zevin et al., 2021; Dall’Amico et al., 2023).
The problem of misinterpreting eccentricity as a potential GR violation is currently a two-fold problem. First, of missing physics; namely, the inclusion of both eccentricity, argument of periapsis (although see Ramos-Buades et al., 2023b), and precession in an inspiral-merger-ringdown waveform model. Distinguishing eccentricity from precession without waveforms that include both (Romero-Shaw et al., 2023) introduces systematic biases in the estimated binary parameters (Romero-Shaw et al., 2020; O’Shea & Kumar, 2021; Favata et al., 2022; Divyajyoti et al., 2024; Dutta Roy & Saini, 2024) that could be misconstrued as false violations of GR (Saini et al., 2022; Bhat et al., 2023; Narayan et al., 2023; Saini et al., 2023; Shaikh et al., 2024). Second, the current analysis methods are producing inconsistent results (Romero-Shaw et al., 2020; Gayathri et al., 2022; Romero-Shaw et al., 2021; O’Shea & Kumar, 2021; Ramos-Buades et al., 2023b; Gupte et al., 2024) for the same events such as GW190521 (Abbott et al., 2020c).
@@ 3.1.2 Tidal Effects
Neutron stars and their mergers are characterized not only by strong gravity but also by extreme matter conditions. To explore how matter affects the space-time deformations around these stars, we need to understand the relation between the dynamical properties of matter and the behavior of strong gravity. Analytic methods are used to model the early inspiral phase of a neutron star binary merger, where neutron stars are approximated as massive point particles with small corrections due to finite-size effects (Flanagan & Hinderer, 2008; Vines et al., 2011; Henry et al., 2020). However, close to merger finite size effects become significant and numerical relativity (NR) simulations are required to capture them accurately (Bernuzzi et al., 2012; Favata, 2014; Wade et al., 2014; Dietrich et al., 2019). Effective one body models achieve a nonperturbative re-summation of the PN information on tidal effects into a complete framework (Bini et al., 2012; Bernuzzi et al., 2012; Bini & Damour, 2014; Bernuzzi et al., 2015; Akcay et al., 2019; Steinhoff et al., 2016; Hinderer et al., 2016); some reduced-order-model versions incorporate NR-calibrated tidal models (Dietrich et al., 2017, 2019; Abac et al., 2024) as also used in Phenomenological models.
@@
The tidal deformation of bodies is directly proportional to the Riemann tensor and its derivatives, produced primarily by the energy-momentum distribution of the companion (Hinderer, 2008), which becomes the second derivatives of the Newtonian potential for the electric-type quadrupole effect in the Newtonian limit. However, such effects are observable in the GWs only if they produce significant mass and current type multipole deformations of the neutron stars in a binary system. The dominant deformations come from the electric-type, $l=2$ tidal deformation, which imprints primarily in the GW phase evolution. However, it is important to note that these tidal effects are relatively small and become more pronounced as the binary approaches merger. While these effects are subtle, their detection has already provided invaluable insights (Abbott et al., 2017a), and with the advent of more advanced detectors (such as XG), we can look forward to even more precise measurements in the near and far future (Pacilio et al., 2022; Kashyap et al., 2022; Evans et al., 2023; Huxford et al., 2023).
@@
The effects of the tidal field on neutron star matter are studied using observed GWs (Abbott et al., 2019a), however, such results are susceptible to waveform systematics and incomplete modeling of neutron star physics. Samajdar & Dietrich (2018); Gamba et al. (2021); Read (2023) show that the inference of tidal parameters with XG detectors can be significantly affected due to waveform systematics. Not including subdominant tidal effects, such as dynamical tides, which become important in the inspiral regime, can also lead to substantial biases in the estimation of tidal parameters (Hinderer et al., 2016; Steinhoff et al., 2016; Pratten et al., 2022; Williams et al., 2022). Likewise, XG detectors will be sensitive to the octupolar electric and quadrupolar magnetic tidal deformabilities, and not including them in the waveform might bias the measurements (Jiménez Forteza et al., 2018).
@@
Resonant mode excitations may contribute distinct features in the waveform from the tidal effect considered in Hinderer (2008). As the inspiraling orbit passes through the frequency of a certain characteristic mode, the resonant excitation of the mode must be compensated by the loss of the same amount of orbital energy, speeding up the following orbital evolution. The excitation of gravity modes (Lai, 1994; Yu & Weinberg, 2017a, b), the interface mode (Tsang et al., 2012; Pan et al., 2020; Lau & Yagi, 2021) and gravitomagnetic mode (Poisson, 2020; Flanagan & Racine, 2007; Ma et al., 2021; Gupta et al., 2023) have been studied, where for the latter two cases the phase modulation may reach the level of $\mathcal{O}(10^{-2})-\mathcal{O}(10^{-1})$ radians in the frequency band of ground-based detectors. Additionally, effects of spins on dynamical tides (Ho & Lai, 1999; Ma et al., 2020; Lai, 1999; Steinhoff et al., 2021), other spin-tidal couplings (Abdelsalhin et al., 2018; Dietrich et al., 2019), spin-induced multipole effects (Poisson, 1998; Krishnendu et al., 2017; Nagar et al., 2018, 2019), nonlinear tides (Yu et al., 2023), higher-order relativistic corrections, and the GW features of tidal disruption in cases with precessing spins (Kawaguchi et al., 2017) are examples of areas requiring further investigations.
@@
Inaccurate or missing physics in analytical and NR modeling due to thermodynamical transformation of nuclear matter during inspiral and post-merger leads to waveform systematics. Such effects include, but not limited to, viscosity (Cutler & Lindblom, 1987; Jones, 2001; Duez et al., 2004; Camelio et al., 2023), thermal effects (Bauswein et al., 2010; Constantinou et al., 2015; Carbone & Schwenk, 2019; Raithel et al., 2021; Fields et al., 2023), phase transition to hyperon condensates or quark matter and other such transformations  (see, e.g., Lattimer, 2012; Baym et al., 2018; Deliyergiyev et al., 2019; Lee et al., 2021; Kain, 2021; Leung et al., 2022) (also see Section 4.3.2 for discussion of proposed exotic matter that has not been observed but, may have compactness close to black holes). As shown in Ripley et al. (2023b, a); Hegade K. R. et al. (2024), the viscous effect introduces a new dissipative channel that modifies the GW phase at 4PN order and higher. If not included in the modeling, a signal that contains such a 4PN effect could be misidentified with a GR deviation at that PN order (and at neighboring PN orders).
@@
Similar effects during the post-merger evolution are subject to systematic bias which requires emphasis on accurate post-merger waveform model development. Currently, only a few post-merger models exist and can detect such effects only in the XG detectors (Breschi et al., 2022, 2019; Soultanis et al., 2022; Prakash et al., 2021, 2023). There are also sources of bias in parameter estimation that are exclusive to data analysis challenges arising from noise systematics. For a minority of events, multiple overlapping signals and confusion background created by CBC mergers could potentially lead to a bias in tidal deformability as described in Section 2.3.
@@
Additionally, GR predicts relations between the spin-induced quadrupole moment and the (quadrupolar, electric) tidal deformability (Yagi & Yunes, 2013a, b, 2017a; Silva et al., 2021) and between tidal deformabilities of different multipolar order and parity (Yagi, 2014) or between different tidal parameters in gravitational waveforms for binary neutron star mergers (Yagi & Yunes, 2016, 2017b) which are only mildly sensitive to the neutron star equation of state. These relations have been used in GW data analyses to reduce the number of search parameters (Chatziioannou et al., 2018; Abbott et al., 2018c) but small equation-of-state variation in the relations can induce systematic biases. One could, however, use constraints on nuclear physics from neutron star observations available at the time to keep updating and reducing the amount of variation in the relations. For example, such variation has been reduced by 50% after GW170817 and current systematic errors on the tidal deformabilities are subdominant than statistical errors until the A# era (Carson et al., 2019). Another way to reduce systematic biases due to the variation in quasi-universal relations is discussed by Kashyap et al. (2022). Since alternative theories predict different relations, an independent measurement of the quantities in the universal relations can therefore be used as null tests of GR, circumventing potential degeneracy with unknown nuclear physics (Yagi & Yunes, 2013a, b, 2017a; Gupta et al., 2018; Maselli et al., 2017; Berti et al., 2024). While the spin-induced quadrupole moment is expected to be small for neutron stars, the magnetic tidal deformability could be measured by XG detectors (Jiménez Forteza et al., 2018).
@@
Besides testing GR, these relations can be used to disentangle source misidentification (discussed in detail in Section 4.3.2), since each model of exotic compact objects other than neutron stars would display their own quasi-universal relation (Maselli et al., 2017; Berti et al., 2024). Notably, the tidal deformability parameter may carry information about the nuclear equation of state and hence offer a unique tool to distinguish conventional neutron stars from the ones with exotic signatures. Analyzing binary neutron star mergers with exotic matter while using waveforms of conventional neutron star binaries could lead to false indications of GR violations. This needs to be investigated thoroughly, so that this effect could be ruled out or observed.
@@
Assuming that our NR-assisted waveform models are accurate and free of systematic biases including those arising from the unknown equation of state, any deviation from the predictions will be indicative of either GR not being the complete theory of gravity or deviations in the coupling of matter to gravity, a subset of which is the test of the strong equivalence principle (Nordtvedt, 1968; Eardley, 1975; Will & Zaglauer, 1989; Will, 1977; Damour & Esposito-Farese, 1992; Will, 1993). Therefore, only after ruling out the systematic effects arising from these inaccuracies, robust conclusions can be drawn about deviations from GR.

### v2

@@ 2.4 Gaps in the Data
Not much work has gone yet to study the effect of data gaps in XG detectors, but some work already exists for data gaps in space-based detectors, from which we can extrapolate some conclusions. Previous work has shown that data gaps can deteriorate and bias parameter estimation for certain sources [96, 97], in particular when the data gap coincides with the merger phase. In general, we would expect that a data gap during the merger would inhibit our ability to constrain deviations from GR at high PN order, while gaps in the early inspiral will do the same for low (or negative) PN order modifications to GR. In particular, if the data has a gap, but our analysis does not account for it, parameter correlations between non-GR and GR parameters are likely to introduce biases that may lead to a false GR violation. Certain methods, such as Bayesian data augmentation [98], however, can be used to include missing data periods as auxiliary variables when sampling the posterior distribution of model parameters that have shown promise at eliminating biases.
@@ 2.5 Detector Calibration Error
The GW strain data $d$ are not directly recorded by the interferometer. Instead, it is reconstructed from the voltage $v(f)$ measured by photodetectors and a response function $R(f)$ that relates the digital readout and GW strain, i.e., $d(f)=R(f)v(f)$ [99]. The calibration process includes a series of measurements to construct a reference model for the response function [99, 100, 101]. Bias in any step of this process can lead to errors in the measured strain data, and systematic errors in parameter estimation could arise if the calibration error is not taken into account. Vitale et al. [102] investigate the consequences of calibration error in Bayesian inference of source parameters. By comparing the ratio between the calibration systematic errors and the statistical uncertainties, their results show that the parameters that suffer the largest biases are those mostly related to the amplitude of GW signals: on average, calibration systematic errors are of $\mathcal{O}(1/100)$ of the statistical uncertainty for sky location parameters and $\mathcal{O}(1/1000)$ for masses. This is potentially because phase errors are localized in frequency and do not accumulate over the inspiral. It implies that calibration errors could have a minor effect in parameterized tests of GR that modify the phase of waveform. Vitale et al. [102] also conclude that $<20$% of amplitude calibration error or $<10-20^{\circ}$ of phase calibration error should not lead to significant biases for all but the loudest signals in the advanced LIGO era, consistent with [103] and [104]. Furthermore, they report that the calibration systematic error is not strongly correlated with SNR as the calibration affects both noise and signal. However, whether such a level of calibration systematics is tolerable in the XG era where SNR can reach $\mathcal{O}(1000)$ is worth investigating. It is still of great importance to improve calibration techniques along with the high sensitivity in the XG era [105, 44].
@@
It is possible to quantify and mitigate calibration errors in detection and data analysis. The uncertainty of the response function can be indicated by the photon calibrators which apply a known radiation pressure directly on the test masses within the detector [106, 107, 108, 99]. Abbott et al. [109] reported $<10$% calibration uncertainty in the strain amplitude and $<5^{\circ}$ in phase during the first LIGO-Virgo observing run , and in the third observing run these uncertainties were reduced to $<7$% and $<4^{\circ}$, respectively [110]. Note that these are overall uncertainties and systematic errors alone are even smaller. These estimates of calibration uncertainties are used as priors to marginalize uncertainties in GW strains during parameter estimation, which effectively mitigates the calibration error [111, 112]. However, this technique might conceal tiny deviations from GR, since it marginalizes over some level of uncertainties on amplitude and phase. Hence, the effect of calibration errors on tests of GR needs to be studied for current and future GW detectors, so that it can be ruled out (or included) as one of the possible causes for false GR violations.
@@ 3.1 Missing Physics in Waveform Models
The current state-of-the-art waveform models used in tests of GR still lack certain physical effects, such as eccentricity of the binary’s orbit, overtones, and non-linearities in the ringdown phase of the binary merger, etc. Including each of these known physical effects individually is crucial for precision GR tests, but their collective inclusion is essential for unbiased assessments of GR. Here we discuss missing physical effects that could lead to a false GR violation.
@@ 3.1.1 Eccentricity
The eccentricity of a binary’s orbit depends on the formation history of the binary. Binaries formed through isolated formation channels in the galactic field are expected to have negligible eccentricity when observed in the frequency band of ground-based detectors, whereas binaries inside dense stellar environments such as globular clusters and nuclear star clusters might have moderate to high eccentricities when observed by these detectors. In an isolated formation channel [113], the binary goes through various mass transfer episodes between its components, and as the components evolve and undergo supernova explosions, the binary orbit could gain some eccentricity due to supernova kicks. However, due to the emission of gravitational radiation [114, 115] the binary’s orbit shrinks, and the binary sheds away all its eccentricity over the long inspiral, leaving it with negligible eccentricity close to merger [114]. For example, if a binary with an initial orbital eccentricity of $0.2$ emits GWs whose dominant mode has a frequency of $0.1$ Hz, the eccentricity reduces to $\sim 10^{-3}$ when it reaches a dominant mode GW frequency of $10$ Hz. That is why binaries detected by LIGO/Virgo are expected to be quasi-circular. On the other hand, a fraction of dynamically formed binaries can still have some eccentricity (and as high as $\sim 1$ at $10$ Hz) when observed in the frequency band of the LIGO/Virgo detectors [116, 117, 118, 119, 120, 121, 122, 123, 124]. Further, environmental effects such as accretion and dynamical friction can also increase the eccentricity of binaries [125].
The problem of misinterpreting eccentricity as a potential GR violation is currently a two-fold problem. First, of missing physics; namely, the inclusion of both eccentricity, argument of periapsis (although see [126]), and precession in an inspiral-merger-ringdown waveform model. Distinguishing eccentricity from precession without waveforms that include both [127] introduces systematic biases in the estimated binary parameters [128, 129, 130, 131, 132] that could be misconstrued as false violations of GR [133, 134, 135, 136, 137]. Second, the current analysis methods are producing inconsistent results [128, 138, 139, 129, 126, 140] for the same events such as GW190521 [141].
@@ 3.1.2 Tidal Effects
Neutron stars and their mergers are characterized not only by strong gravity but also by extreme matter conditions. To explore how matter affects the space-time deformations around these stars, we need to understand the relation between the dynamical properties of matter and the behavior of strong gravity. Analytic methods are used to model the early inspiral phase of a neutron star binary merger, where neutron stars are approximated as massive point particles with small corrections due to finite-size effects [142, 143, 144]. However, close to the merger finite size effects become significant and numerical relativity (NR) simulations are required to capture them accurately [145, 146, 147, 148]. Effective one body models achieve a nonperturbative re-summation of the PN information on tidal effects into a complete framework [149, 145, 150, 151, 152, 153, 154]; some reduced-order-model versions incorporate NR-calibrated tidal models [155, 148, 156] as also used in Phenomenological models.
@@
The tidal deformation of bodies is directly proportional to the Riemann tensor and its derivatives, produced primarily by the energy-momentum distribution of the companion [157], which becomes the second derivatives of the Newtonian potential for the electric-type quadrupole effect in the Newtonian limit. However, such effects are observable in the GWs only if they produce significant mass and current type multipole deformations of the neutron stars in a binary system. The dominant deformations come from the electric-type, $l=2$ tidal deformation, which imprints primarily in the GW phase evolution. However, it is important to note that these tidal effects are relatively small and become more pronounced as the binary approaches merger. While these effects are subtle, their detection has already provided invaluable insights [62], and with the advent of more advanced detectors (such as XG), we can look forward to even more precise measurements in the near and far future [158, 159, 160, 161].
@@
The effects of the tidal field on neutron star matter are studied using observed GWs [2], however, such results are susceptible to waveform systematics and incomplete modeling of neutron star physics. For example, Refs. [162, 163, 164] show that the inference of tidal parameters with XG detectors can be significantly affected due to waveform systematics. Not including subdominant tidal effects, such as dynamical tides, which become important in the inspiral regime, can also lead to substantial biases in the estimation of tidal parameters [154, 153, 165, 166]. Likewise, the effects of spins on dynamical tides [167, 168, 169, 170], other spin-tidal couplings [171, 148], spin-induced multipole effects [172, 173, 174, 175], nonlinear tides [176], higher-order relativistic corrections, and the GW features of tidal disruption in cases with precessing spins [177] are examples of areas requiring further investigations. Further, XG detectors will be sensitive to the octupolar electric and quadrupolar magnetic tidal deformabilities, and not including them in the waveform might bias the measurements [178].
@@
Resonant mode excitations may contribute distinct features in the waveform from the tidal effect considered in [157]. As the inspiraling orbit passes through the frequency of a certain characteristic mode, the resonant excitation of the mode must be compensated by the loss of the same amount of orbital energy, speeding up the following orbital evolution. The excitation of gravity modes [179, 180, 181], the interface mode [182, 183, 184] and gravitomagnetic mode [185, 186, 187, 188] have been studied, where for the latter two cases the phase modulation may reach the level of $\mathcal{O}(10^{-2})-\mathcal{O}(10^{-1})$ radians in the frequency band of ground-based detectors.
@@
Inaccurate or missing physics in analytical and NR modeling due to thermodynamical transformation of nuclear matter during inspiral and post-merger leads to waveform systematics. Such effects include, but not limited to, viscosity [189, 190, 191, 192], thermal effects [193, 194, 195, 196, 197], phase transition to hyperon condensates or quark matter and other such transformations (see, e.g., [198, 199, 200, 201, 202, 203] and also see Section 4.3.2 for discussion of proposed exotic matter that has not been observed but, may have compactness close to black holes). As shown in [204, 205, 206], the viscous effect introduces a new dissipative channel that modifies the GW phase at 4PN order and higher. If not included in the modeling, a signal containing such a 4PN effect could be misinterpreted as a GR deviation at that PN order and at neighboring PN orders.
@@
Similar effects during the post-merger evolution are subject to systematic bias which requires emphasis on accurate post-merger waveform model development. Currently, only a few post-merger models exist and can detect such effects only in the XG detectors [207, 208, 209, 210, 211]. There are also sources of bias in parameter estimation that are exclusive to data analysis challenges arising from noise systematics. For a minority of events, multiple overlapping signals and confusion background created by CBC mergers could potentially lead to a bias in tidal deformability as described in Section 2.3.
@@
Additionally, GR predicts relations between the spin-induced quadrupole moment and the (quadrupolar, electric) tidal deformability [212, 213, 214, 8] and between tidal deformabilities of different multipolar order and parity [215] or between different tidal parameters in gravitational waveforms for binary neutron star mergers [216, 217] which are only mildly sensitive to the neutron star equation of state. These relations have been used in GW data analyses to reduce the number of search parameters [218, 219] but small equation-of-state variation in these relations can induce systematic biases. One could, however, use constraints on nuclear physics from neutron star observations available at the time to keep updating and reducing the amount of variation in the relations. For example, such variation has been reduced by 50% after GW170817 and current systematic errors on the tidal deformabilities are subdominant than statistical errors until the A# era [220]. Another way to reduce systematic biases due to the variation in quasi-universal relations is discussed by [159]. It should be noted that the subpercent accuracy in the universal relations will become important in deducing the correct equation of state and hence in the tests of GR for the sensitivities corresponding to XG detectors. Since alternative theories predict different relations, an independent measurement of the quantities in the universal relations can therefore be used as null tests of GR, circumventing potential degeneracy with unknown nuclear physics [212, 213, 214, 221, 222, 223]. While the spin-induced quadrupole moment is expected to be small for neutron stars, the magnetic tidal deformability could be measured by XG detectors [178] and might need to be included in the waveform models.
@@
Besides testing GR, these relations can be used to disentangle source misidentification (discussed in detail in Section 4.3.2), since each model of exotic compact objects other than neutron stars would display their own quasi-universal relation [222, 223]. Notably, the tidal deformability parameter may carry information about the nuclear equation of state and hence offer a unique tool to distinguish conventional neutron stars from the ones with exotic signatures. Analyzing binary neutron star mergers with exotic matter while using waveforms of conventional neutron star binaries could lead to false indications of GR violations. This needs to be investigated thoroughly, so that this effect could be ruled out or observed.
@@
Assuming that our NR-assisted waveform models are accurate and free of systematic biases including those arising from the unknown equation of state, any deviation from the predictions will be indicative of either GR not being the complete theory of gravity or deviations in the coupling of matter to gravity, a subset of which is the test of the strong equivalence principle [224, 225, 226, 227, 228, 229]. Therefore, only after ruling out the systematic effects arising from these inaccuracies, robust conclusions can be drawn about deviations from GR.